OpenStudy (anonymous):

What function is quadratic? a) y=x(3x+2) b) x-5y=9 c)y=5x-2 d)y=x3+2x2

OpenStudy (anonymous):

i dont know

OpenStudy (solomonzelman):

quadratic function is in the form of \(\large\color{black}{ \displaystyle y=\color{red}{ a}{\rm x}^2+\color{green}{ b}{\rm x}+ \color{blue}{ c} }\) Where \(\large\color{black}{ \displaystyle \color{red}{ a} }\) , \(\large\color{black}{ \displaystyle \color{green}{ b} }\) and \(\large\color{black}{ \displaystyle \color{blue}{ c} }\) are just numbers. \(\large\color{black}{ \displaystyle \color{green}{ b} }\) and \(\large\color{black}{ \displaystyle \color{blue}{ c} }\) could be any positive/negative number or zero. And, \(\large\color{black}{ \displaystyle \color{red}{ a}\ne0 }\)

OpenStudy (solomonzelman):

expand all of the choices, and see which one is like \(\large\color{black}{ \displaystyle \large\color{black}{ \displaystyle y=\color{red}{ a}{\rm x}^2~~... } }\)